OpenStudy (anonymous):

Linear Function Problem. A swimming pool contains 1500 gallons of water. The pool is being drained for the season @ a rate of 35gallons per minute. Find a rule for the linear function that describes the amount of water in the tank. Use it to determine how much will be in the tank after 25 minutes.

OpenStudy (anonymous):

total water = 1500 total water after a minute = 1500 - 35 total water after t minutes = 1500 - 35 * t w = 1500 - 35 t w (25) = 1500 - 35 * 25 = 1500 - 875 = 625

OpenStudy (anonymous):

... woah. I still don't get the rule for the linear function though 0.o

OpenStudy (anonymous):

You have a total of 1,500 gallons, if each minute you drain 35 gallons, after each minute there's 35 gallons less. After one minute you drained 35 gallons, there were 1,500 at the beginning, how much is there now? 1,500 - 35 = 1,465, after two minutes 1,500 - 70 = 1,430, and so on.

OpenStudy (anonymous):

Linear function means y = x to the first and whatever else. In this case, t is to the first power. f(t) = 1500 - 35t will give you the answer, where t is time.

OpenStudy (anonymous):

ohh okai. thank you